OpenStudy (anonymous):

solve using the addition and multiplication principles..1+3x<28

OpenStudy (angela210793):

1+3x<28(-1 both sides) 3x<28-1 3x<27(divide with 3 both parts) x<27/3 x<9 x belongs to]-infinite;9[

OpenStudy (anonymous):

\[\left\{ x \left| x \ge9 \right| \right\}\]

OpenStudy (anonymous):

that is the answer I have

OpenStudy (anonymous):

1+3x<28(-1 both sides) 3x<27(divide both sides with 3) x<9

OpenStudy (angela210793):

well no...the answer is the one i found....y did u equal it?

OpenStudy (anonymous):

i have to put it into a solution set. oh i clicked on the wrong sign

OpenStudy (angela210793):

w8

OpenStudy (anonymous):

it should be that shouldnt it?

OpenStudy (angela210793):

\[x \in]-\infty;9[\]

OpenStudy (angela210793):

x must be only less than 9, not 9 or less

OpenStudy (anonymous):

if i put 9 in for x it would be 1+3(9)<28 1+28<28 29<28 that look about right?

OpenStudy (anonymous):

nope thats not right either

OpenStudy (angela210793):

so the inequality isn't right cause 29>28

OpenStudy (anonymous):

1+3(9)<28 1+27<28 28=28

OpenStudy (angela210793):

anyway 9 would be included only if it said \[1+3x \le28\]

OpenStudy (anonymous):

so i just need to change my symbol to the other direction?

OpenStudy (angela210793):

the symbol must be <

OpenStudy (angela210793):

Got it Tena?

OpenStudy (anonymous):

yes thanks.

OpenStudy (angela210793):

:)

OpenStudy (anonymous):

good angela

OpenStudy (angela210793):

:P:P:P I did pretty good in my math exam :):)

OpenStudy (anonymous):

heyy yea congrats

OpenStudy (angela210793):

Thnx :)